Chapter 2 Homework M MHE Reader x M MHE Reader ework Help Se Saved You are an assistant to a senator who chairs an ad hoc committee on reforming taxes on telecommunication services Based on research, AT&T has spent over $15 million on related paperwork and compliance costs. Moreover, depending on the locale, telecom taxes can amount to as much as 25 percent of a consumers phone bill. These high tax rates on telecom services have become quit controversial, due to the fact that the deregulation of the telecom industry has led to a highly competitive market. Your best indicate that, based on current tax rates, the monthly market demand for telecommunication services is given by Q-300-4P the market supply (including taxes) is QS - 3P- 120 (both in millions, where P is the monthly price of the of The senator is considering tax reform that 3.2P-120. How much money per unit would a typical consumer save each month as a resuit of the proposed legislation Instruction: Enter your response rounded to the nearest penny two decimal be Prev13 of 18 E Score answer > 0
Answer #1

In pre-tax equilibrium, Qd = Qs.

300 - 4P = 3P - 120

7P = 420

P = 60

Q = 200 - (4 x 60) = 300 - 240 = 60 (million)

In after-tax equilibrium, Qd = Qs.

300 - 4P = 3.2P - 120

7.2P = 420

P = 58.33 (Price paid by buyers)

Saving per unit = Decrease in price paid by buyers = 60 - 58.33 = 1.67